In a heat pump, heat from the outdoors at –5C is transferred to a room at 17C, energy being supplied by an electric motor. How many joules of heat will be delivered to the room for each joule of electric energy consumed, ideally?
1
Expert's answer
2019-12-16T11:11:43-0500
"\frac{Q_h}{T_h}=\frac{Q_l}{T_l}"

From the conservation of energy:

"W+Q_l=Q_h"

"Q_h=(Q_h-W)\frac{T_h}{T_l}"

"Q_h=(Q_h-1)\frac{290.15}{268.15}"

"Q_h=13\ J"